# On the emergence of numerical instabilities in Next Generation Reservoir Computing

### v1

@@ 5.3. Dependence on the length of training data
As we have seen, the statistical properties of the original dynamics give insights into the condition number of $\Psi$. This is not different from the length of training data $N_{\mathrm{train}}$. As aforementioned, to define the least squares problem during training, $\Psi$ should be tall and slim ($N_{\mathrm{train}}>m$); otherwise, the matrix has a non-trivial kernel and is rank deficient. To test the dependence on $N_{\mathrm{train}}$, Fig. 5 shows the condition number of $\Psi$ for $N_{\mathrm{train}}>\binom{kd+p}{p}$. For moderate values of $N_{\mathrm{train}}$, that is not much larger than $\binom{kd+p}{p}$, the condition number is not small. We observe that the condition number decays monotonically towards a limiting value $10^{2}$, but not uniformly. Initially, the decay is much faster than compared to larger values of $N_{\mathrm{train}}$. In the left panel, the NGRC model corresponds to the reference model we have been using so far, $k=1$ and $p=2$, and the right panel corresponds to the NGRC model with more terms, $k=2$, $\tau=50$, and $p=2$.
@@
Since we are dealing with an ordinary differential equation, the length of the training data is determined by the time step $h$. Both panels of Fig. 5 show the dependence of $\kappa(\hat{\Psi})$ for different time steps. For this experiment, the numerical integration is performed using the explicit forward Euler method with time step $h=0.01$, then the time series is subsampled accordingly to the larger time steps. Similarly to $\tau$, the convergence to the limiting condition number is faster as the time step $h$ increases.
@@
Notice that in both cases the condition number decays, but as expected, an NGRC model that contains more monomials requires more data points to reach the limit. In other words, the linear dependencies among columns can be long-lived as more monomials are used in the NGRC model. The convergence of $\kappa(\Psi)$ towards the limit and the difference between small and large $h$ can be explained due to the statistical properties of the Lorenz system, as we presented. The convergence follows from the Birkhoff averaging theorem in Eq. 18. For sufficiently long training data, the entries of $\hat{\Psi}^{\top}\Psi$ converge in a neighborhood of a limit value under different rates that depend on the pair of observables $(\psi_{i}\psi_{j})\circ\mathbf{g}_{k,\tau}$. This is in agreement with a recent result in traditional reservoir computing, which has shown that the Tikhonov regularization is an $L^{2}(\mu)$ approximation of ergodic dynamics [HHD21].
@@
Similarly, $h$ in this particular construction plays a similar role as $\tau$ before. Larger $h$ corresponds to skipping the time evaluation of the monomial along time, and the exponential decay of correlations results in a smaller condition number. The fact that the entries of the matrix converge to a limit motivates scaling the matrix by the factor $\frac{1}{\sqrt{N_{\mathrm{train}}}}$. This is also relevant for the regularizer parameter, as we detail.
@@
Remark 5.2 (Scaled regularizer parameter $\beta$ by the length of training data $N_{\mathrm{train}}$).
@@
The scaling $\frac{1}{\sqrt{N_{\mathrm{train}}}}$ in the definition of $\Psi$ in Eq. 5 also scales $\beta$ with respect to the data length $N_{\mathrm{train}}$, which has been a common practice in statistics [GHW79], and more recently in reservoir computing [CAAG24, ZdSC25]. Let us denote $\tilde{\Psi}=\frac{1}{\sqrt{N_{\mathrm{train}}}}\Psi$. It follows that when $\tilde{\Psi}$ is used, the coefficient $\mathbf{w}(\beta)$ in Eq. 7 can be mapped to another coefficient with an adjusted regularizer parameter $N_{\mathrm{train}}\beta$. Note that
@@
6. Dynamical instability is algorithm-dependent
Figure 6. Ill-conditioning leads to numerical algorithm dependency. Comparison of three different numerical algorithms: Cholesky, SVD, and LU as the regularizer parameter $\beta$ is increased. Box plots in top to bottom panels as follows: valid prediction time (VPT), the distance between the induced successive local maxima maps, and the error $E$ between the power spectrum density, respectively. The colored hashed areas correspond to: absent/small (green), large (purple) regularization, and level of regularization in which all algorithms perform similarly (orange). The error bars are with respect to 50 different initial conditions. The orange band corresponds to the regularizer value in which all algorithms perform equally within statistical confidence. Whenever the algorithm failed to find a solution, we set $\mathbf{w}_{i}(\beta)=0$. Moreover, if the trajectory of the NGRC during testing was unbounded, we set the metric value to $-1$. This explains the blank spaces for Cholesky and LU decompositions. The parameters are $h=0.01$, delay dimension $k=2$, time lag $\tau=1$, the maximum degree $p=2$, and $N_{\mathrm{train}}=5000$.
As we have seen in Section 5, $\Psi$ can be ill-conditioned depending on the choices of maximum degree of polynomial, delay dimension, time lag, and length of training data. In this section, we consider the ill-conditioned regime for small $\tau$ and $h$, as commonly done in NGRC numerical experiments [GBGB21, GFR22, ZC23]. While Section 3.5 analyzed the algorithm’s accuracy under perturbations during NGRC training, here we focus on how such differences manifest in the NGRC’s resulting dynamics
This test mirrors a common practice in data-driven machine learning: fixing the hyperparameters, increasing the regularization parameter, and observing performance. This experiment highlights how different algorithms for solving the least-squares problem behave under varying regularization levels. We consider the NGRC model with $k=2$, $\tau=1$, $p=2$, and varying $\beta$ values, see horizontal axis of Fig. 6. The NGRC performance during testing is quantified by topological and statistical metrics, as described in detail in Section 8.1: valid prediction time (VPT) [VPH+20] quantifies for how long, in Lyapunov times, the NGRC trajectory is close to the original one; the distance between the induced successive local maxima map; and divergence error $E$ between the power spectrum densities estimated from the trajectories. Note that the columns of $\Psi$ are not normalized, as normalization can lead to different solutions [GVL96].
Fig. 6 shows the NGRC performance across different values of $\beta$ using the three algorithms. Starting from the right, the purple region corresponds to the regime of large $\beta$. The VPT decreases monotonically, where all three algorithms obtain a solution whose performance drops because the regularization is too strong. The distance between successive maps and the divergence error remain statistically indistinguishable across initial conditions, until they deteriorate due to excessive regularization.
The orange region corresponds to the regularizer value that all algorithms perform equally over all metrics. This regime corresponds to sufficient regularization for best performance in all three metrics. As opposed to smaller $\beta$, where the algorithms should overfit the input data, being susceptible to the influence of the ill-conditioning of $\Psi$. But this is not the case. The green regions show that for smaller $\beta$, the three algorithms yield different NGRC performances. Cholesky and LU decompositions exhibit discontinuous behavior as $\beta$ changes, whereas SVD varies smoothly. SVD is robust against the ill-conditioning of $\Psi$. Surprisingly, at $\beta=0$, its NGRC performance remains unaltered. Thus, regularization becomes unnecessary when an appropriate algorithm, such as SVD, is used. This result may have gone unnoticed, as the common practice is to apply sufficiently strong regularization to ensure good NGRC performance. And naturally, it points out to: we could attempt to minimize the condition number to avoid any conditioning-induced instabilities. In the next section, we show that minimizing the condition number does not guarantee dynamical stability.
6.1. Minimizing condition number does not imply dynamical stability
Figure 7. Better conditioning does not imply higher dynamical stability. The dots correspond to the median with respect to 25 different initial conditions, and error bars are $25\%$ (lower bound) and $75\%$ (upper bound) quantiles. The orange band represents the $\tau=15$, which shows the NGRC model with best performance in both metrics, VPT and divergence error $E$, but condition number $\kappa(\Psi)$ at $10^{4}$, which is not the minimum condition number possible with respect to $\tau$. If the trajectory of the NGRC during testing was unbounded, we set the metric value to $-1$. The parameters are $h=0.01$, delay dimension $k=2$, the maximum degree $p=5$, $N_{\mathrm{train}}=10000$ and regularizer parameter $\beta=0$.
In this section, we discuss an important observation: better conditioning of the feature matrix does not guarantee dynamical stability. For this test, we consider a partial measurement of the Lorenz system, having access only to the $x-$component time series $\{x_{n}\}_{n\geq 0}$, which is normalized to be in $[-1,1]$ as previous sections, see the left panel of Fig. 8. Recently, it has been shown that NGRC can forecast the $x-$component time series [RP24]. The underlying dynamics is not the Euler discretization of a polynomial vector field, but an unknown function of the current time and time-delayed coordinates by Takens’ embedding theorem [PCFS80, Tak81]. Thus, the unknown function does not necessarily lie in the span of $\mathcal{P}_{p}^{kd}$ and must be approximated by them, a scenario more typical in NGRC applications, and distinct from earlier sections.
The NGRC model has a maximum degree $p=5$ and $k=3$, which is chosen based on the standard false nearest neighbors method for estimating embedding delay dimension [KS03]. For the time lag, we proceed as follows: we increase the $\tau$ and measure the NGRC performance by the valid prediction time and divergence error between the power spectrum densities, see Fig. 7.
As expected by our analysis in Section 5.2, increasing $\tau$ reduces the condition number monotonically. However, NGRC performance exhibits a different trend. At first, the NGRC performance has a peak at $\tau=15$ highlighted by the orange band. But then, the performance worsens. This agrees with former observations in constructing embeddings of partial measurements of chaotic systems. For small $\tau$, the $x-$component time series is highly correlated with each other, while for large $\tau$, the time series become highly uncorrelated. In both cases, capturing the embedding becomes challenging [FS86, KS03]. The good time lag $\tau$ corresponds to the first minimum of mutual information. Interestingly, the NGRC performance has a peak and coincides with the time lag that attains the first minimum of mutual information. This illustrates that although better conditioning ensures numerical stability during training, it does not guarantee dynamical stability in NGRC predictions. An example of the NGRC trajectory and power spectrum at $\tau=15$ is depicted in Fig. 8.
@@ 6.1. Minimizing condition number does not imply dynamical stability
Figure 8. NGRC learns partial measurements of the Lorenz system. The left panel shows the NGRC trajectory (red) compared to the original (blue). The right panel shows the comparison between the power spectrum densities. NGRC The parameters are $h=0.01$, delay dimension $k=2$, time lag $\tau=15$, the maximum degree $p=5$, $N_{\mathrm{train}}=10000$ and regularizer parameter $\beta=0$.
7. Discussion and conclusions
This work has presented a detailed numerical analysis of the feature matrix conditioning in NGRC. By leveraging the Vandermonde- and Hankel-like structures of the feature matrix and the statistical properties of the Lorenz system, we elucidated the interplay between hyperparameters and feature matrix conditioning. Through numerical experiments, we demonstrated that the condition number of $\Psi$ can become prohibitively large: (i) high maximum degree $p$, (ii) short time delay $\tau$, and (iii) moderate length of training data $N_{\mathrm{train}}$. We also observed that the choice of numerical algorithm used for training substantially impacts robustness to ill-conditioning.
Incidentally, in the presence of an ill-conditioned feature matrix, SVD-based training can lead to a stable NGRC even without regularization. This dependence on the numerical method suggests a practical step: evaluate the conditioning of the feature matrix and prefer SVD-based training before deciding whether regularization is necessary. While SVD may be computationally expensive for large systems, alternatives for efficient Tikhonov regularization exist, including classical approaches such as Élden’s method [Eld77] and modern techniques for large-scale inverse problems [LRV12].
Interestingly, while bad conditioning may cause NGRC instability, it does not imply it. For instance, we showed that a successful reconstruction of the $x$-component dynamics remains possible despite a non-minimal condition number of $\Psi$. This correspondence of conditioning and NGRC stability suggests that minimizing the condition number is not a universally optimal objective. In fact, classical numerical analysis states that conditioning is highly sensitive to the distribution of points being evaluated (collocation points or nodes) and the choice of polynomial functions [KRS19]. For example, Vandermonde matrices evaluated at points nearly equally spaced on the unit circle or using orthogonal polynomial bases (like Chebyshev) are known to be well-conditioned. These properties, however, are unrelated to the underlying dynamics. This observation raises an exciting research direction of constructing a basis of orthogonal polynomials adapted to the distribution of points generated by the trajectory of the dynamical system. A first step towards this direction has been applied to coupled maps for network reconstruction [PdSvS23].

### v2

@@ IV.3 Dependence on the length of training data
As we have seen, the statistical properties of the original dynamics give insights into the condition number of $\Psi$. This is not different from the length of training data $N_{\mathrm{train}}$. As aforementioned, to define the least squares problem during training, $\Psi$ should be tall and slim ($N_{\mathrm{train}}>m$); otherwise, the matrix has a non-trivial kernel and is rank deficient. To test the dependence on $N_{\mathrm{train}}$, Fig.˜7 shows the condition number of $\Psi$ for $N_{\mathrm{train}}>\binom{kd+p}{p}$. For moderate values of $N_{\mathrm{train}}$, which is not much larger than $\binom{kd+p}{p}$, the condition number is not small. We observe that the condition number decays monotonically toward a limiting value $10^{2}$, but not uniformly. Initially, the decay is much faster than compared to larger values of $N_{\mathrm{train}}$. In the left panel, the NGRC model corresponds to the reference model we have been using so far, $k=1$ and $p=2$, and the right panel corresponds to the NGRC model with more terms, $k=2$, $\tau=50$, and $p=2$.
@@
Since we are dealing with an ordinary differential equation, the length of the training data is determined by the time step $h$. Both panels of Fig.˜7 show the dependence $\kappa(\hat{\Psi}) fo$r different time steps. For this experiment, the numerical integration is performed using the explicit forward Euler method with time step $h=0.01$, then the time series is subsampled accordingly to the larger time steps. Similarly to $\tau$, the convergence to the limiting condition number is faster as the time step $h$ increases.
@@
Notice that in both cases the condition number decays, but as expected, an NGRC model that contains more monomials requires more data points to reach the limit. In other words, the linear dependencies among columns can be long-lived as more monomials are used in the NGRC model. The convergence of $\kappa(\Psi)$ toward the limit and the difference between small and large $h$ can be explained due to the statistical properties of the Lorenz system, as we presented. The convergence follows from the Birkhoff averaging theorem in Eq.˜19. For sufficiently long training data, the entries of $\hat{\Psi}^{\top}\Psi$ converge in a neighborhood of a limit value under different rates that depend on the pair of observables $(\psi_{i}\psi_{j})\circ\mathbf{g}_{k,\tau}$. This is in agreement with a recent result in traditional reservoir computing, which has shown that the Tikhonov regularization is an $L^{2}(\mu)$ approximation of ergodic dynamics [55].
@@
Similarly, $h$ in this particular construction plays a similar role as $\tau$ before. Larger $h$ corresponds to skipping the time evaluation of the monomial along time, and the exponential decay of correlations results in a smaller condition number. The fact that the entries of the matrix converge to a limit motivates scaling the matrix by the factor $\frac{1}{\sqrt{N_{\mathrm{train}}}}$. This is also relevant for the regularizer parameter, as we detail in the following remark.
@@
Remark IV.3 (Scaled regularizer parameter $\beta$ by the length of training data $N_{\mathrm{train}}$).
@@
The scaling $\frac{1}{\sqrt{N_{\mathrm{train}}}}$ in the definition of $\Psi$ in Eq.˜5 also scales $\beta$ with respect to the data length $N_{\mathrm{train}}$, which has been a common practice in statistics [56], and more recently in reservoir computing [57, 26]. Let us denote $\tilde{\Psi}=\frac{1}{\sqrt{N_{\mathrm{train}}}}\Psi$. It follows that when $\tilde{\Psi}$ is used, the coefficient $\mathbf{w}(\beta)$ in Eq.˜7 can be mapped to another coefficient with an adjusted regularizer parameter $N_{\mathrm{train}}\beta$. Note that
@@
V Condition number guides better choices of hyperparameter
The characterization of the condition number from Section˜IV can be used to guide better choices of hyperparameter for NGRC training. In this section, we evaluate NGRC in two application scenarios: (i) sparsely-sampled time series; and (ii) partial measurement, where only the $x$-coordinate of the Lorenz system is available. In both cases, we identify regions of the hyperparameter space where the NGRC model reproduces the original dynamics without any regularization, thereby avoiding the extra computational cost of tuning $\beta$.
To test robustness against higher-accuracy numerical integrators, throughout this section, we utilize a fourth-order Runge–Kutta method with a fine time step of $0.001$, followed by subsampling by time step $h$. Also, the length of training data is sufficient to avoid ill-conditioning due to undersampling, as discussed in Section˜IV.3. Additional numerical experiments with the electronic circuit model of the Double-Scroll introduced in [58] generalize the findings to other dynamical systems, see Supplementary Material.
V.1 Sparsely-sampled training data
Figure 8: Sparsely-sampled time series requires an NGRC model with a larger polynomial degree. The top row shows the maximum closeness of fit (black) and the condition number of $\Psi$ (red) across polynomial degrees, for the three numerical algorithms. Dots indicate the median over $25$ initial conditions. Subsequent rows show NGRC performance metrics as box plots over the different initial conditions and bounded models, similar to Fig.˜4. Blank spaces correspond to unbounded NGRC models over all initial conditions. Training data were generated with a fourth-order Runge–Kutta and subsampled with step size $h=0.1$. The parameters are delay dimension $k=1$, time lag $\tau=1$, $N_{\mathrm{train}}=1000$, $N_{\mathrm{test}}=1000$, and $\beta=0$.
Consider the scenario where the Lorenz system is subsampled with a moderate time step $h=0.1$. The input data exhibit more intricate dynamics than the original quadratic vector field. The effective dynamics result from a combination of the numerical approximation introduced by the numerical integrator and successive polynomial compositions between time samples. Consequently, capturing this behavior requires a higher-degree polynomial expansion so that the true dynamics lie within the span of $\mathcal{P}_{p}^{kd}$. In this context, we evaluate the NGRC performance as the maximum degree $p$ is increased, see Fig.˜8.
Increasing $p$ to match the original dynamics inevitably pushes the NGRC model toward an ill-conditioned regime, as analyzed in Section˜IV.1. To examine this, we also test the consistency of the numerical algorithms’ behavior discussed in Section˜II.7. The top row of Fig.˜8 displays the trade-off between the maximum closeness of fit $\theta_{\max}$ and the condition number as $p$ increases. The condition number grows exponentially, consistent with the same growth rate found for the explicit Euler method in Section˜IV.1. Unlike increasing $\beta$ in Fig.˜3, here the fitting accuracy improves with larger $p$, since a richer polynomial basis better represents the dynamics. Only the SVD algorithm captures this improvement smoothly, showing a monotonic increase in fitting accuracy. By contrast, the Cholesky and LU solutions degrade rapidly beyond $p>7$ due to numerical instability.
To quantify the NGRC performance during the testing phase, we evaluate the valid prediction time, the distance between the induced successive local maxima maps, and the divergence error $E$. The main result is that, for certain values of $p$, NGRC can successfully reproduce the input data without any regularization. Moreover, examining how different numerical algorithms perform, Cholesky and SVD exhibit statistically similar performance within the orange shaded area, whereas the LU algorithm fails for all values of $p$. Interestingly, similar to the dependence on $\beta$, the SVD algorithm displays a smooth and stable dependence on $p$. It remains accurate across all tested values, with performance comparable to the orange shaded region even as $p$ increases. In contrast, Cholesky breaks down for $p>7$, producing different dynamics from the original system. This demonstrates that the SVD algorithm is consistently more robust to the ill-conditioning of the feature matrix, making it a better choice when aiming for an accurate NGRC model without regularization.
V.2 Partial measurement from the Lorenz system
@@
Figure 9: NGRC learns partial measurements of the Lorenz system. The left panel shows the NGRC trajectory (red) compared to the original (blue). The right panel shows the comparison between the power spectrum densities. The numerical integration is the fourth-order Runge-Kutta with time step $h=0.01$. The parameters are delay dimension $k=3$, time lag $\tau=15$, the maximum degree $p=7$, $N_{\mathrm{train}}=10000$, $N_{\mathrm{test}}=10000$ and regularizer parameter $\beta=0$.
Consider a partial measurement of the Lorenz system, where only the $x$-coordinate time series $\{x_{n}\}_{n\geq 0}$ is available, as shown in the left panel of Fig.˜9. According to Takens’ embedding theorem [59, 60], the underlying dynamics in this case is an unknown function of the current and time-delayed coordinates, rather than the Euler discretization of a polynomial vector field. Consequently, this function does not necessarily lie in the span of $\mathcal{P}_{p}^{kd}$ and must instead be approximated.
Rather than performing a grid search over all hyperparameters, we analyze NGRC performance by varying only the time lag $\tau$, while keeping the delay dimension $k$ and maximum polynomial degree $p$ fixed. The delay dimension $k=3$ is chosen using the false nearest neighbors method [4], and the polynomial degree is set high enough to approximate the underlying dynamics. To demonstrate robustness, we test the forecasting task on data generated by two numerical integrators for $h=0.01$: the explicit Euler method (as in earlier sections) and the fourth-order Runge–Kutta method, which accounts for subsampling. All simulations use the SVD algorithm, as it provides the most stable numerical algorithm.
Fig.˜10 shows the closeness of fit $\theta_{x}$ and the condition number $\kappa(\Psi)$ as $\tau$ increases. For both numerical integrators, increasing $\tau$ leads to a monotonic increase in $\theta_{x}$ and a monotonic decrease in $\kappa(\Psi)$, consistent with our earlier discussion in Section˜IV.2. In particular, for small $\tau$, the columns of the feature matrix are highly correlated, but as $\tau$ grows, they become increasingly uncorrelated. While this improves numerical stability, it reduces the NGRC model’s ability to capture the dynamics accurately. This might explain the plateau (visible in logarithmic scale) of $\theta_{x}$ for $\tau>20$.
The NGRC performance is evaluated using the valid prediction time, the divergence error $E$ between the power spectrum densities, and the fraction of bounded models, defined as the percentage of initial conditions whose autonomous NGRC trajectory remains within the compact region $1,1]^{3}$. The key result is that, for both numerical integrators, NGRC can reproduce the partial measurement dynamics without any regularization, see an example trajectory and power spectrum for $\tau=15$ in Fig.˜9. However, the dependence on $\tau$ differs between the two types of data. For the explicit Euler method, a broad range of $\tau$ values yields stable, bounded models with similar performance. In contrast, the Runge–Kutta data exhibit a narrower window of good performance, with only three $\tau$ values achieving a fraction larger than $0.5$ of bounded NGRC models. This difference likely arises from the additional complexity introduced by partial observation combined with subsampling in the Runge–Kutta data.
Interestingly, the optimal NGRC performance for both data types coincides with the time lag $\tau=15$ corresponding to the first minimum of mutual information, a standard heuristic for selecting embedding time lags [61, 4]. This suggests that the method used for embedding time series can inform the choice of NGRC hyperparameters, reducing the need for exhaustive grid searches and accelerating training.
Figure 10: Methods used for embedding time series may inform NGRC modelling of partial measurements. The top row displays the closeness of fit $\theta_{x}$ (black) and condition number of $\Psi$ (red) with respect to the increasing time lag $\tau$. The subsequent rows show the NGRC performance in both metrics, VPT and divergence error $E$. The orange band represents the $\tau=15$, which is informed by the first minimum of mutual information. The blue colormap represents the fraction of bounded NGRC models over the $25$ initial conditions during testing. The dots correspond to the median, and error bars are $25\%$ (lower bound) and $75\%$ (upper bound) quantiles, which are shown only for NGRC models with a fraction larger than $0.5$. The maximum degree is $p=5$ for the explicit Euler method, and $p=7$ for the fourth-order Runge-Kutta method. The parameters are $h=0.01$, delay dimension $k=3$, $N_{\mathrm{train}}=10000$, $N_{\mathrm{test}}=10000$ and regularizer parameter $\beta=0$.
VI Discussion and conclusions
This work demonstrated that numerical analysis of NGRC training can guide better hyperparameter choices, often enabling training without any regularization. We have two main contributions. First, by leveraging the Vandermonde- and Hankel-like structures of the feature matrix and the statistical properties of the chaotic systems, we elucidated the interplay between hyperparameters and feature matrix conditioning. Through numerical experiments, we demonstrated that the condition number of $\Psi$ can become prohibitively large: (i) high maximum degree $p$, (ii) short time delay $\tau$, and (iii) short length of training data $N_{\mathrm{train}}$. Second, we showed that the choice of numerical algorithm critically impacts robustness to ill-conditioning. Across all scenarios tested, the LU algorithm was significantly less stable than the Cholesky and SVD algorithms. Moreover, there exist hyperparameter regions where SVD-based training achieves accurate forecasts without any regularization. This result may have gone unnoticed, as the common practice is to apply sufficiently strong regularization to ensure good NGRC performance, eliminating the large difference between the algorithms’ solutions, as we have seen in Section˜II.7. This dependence on the numerical method suggests a practical step: prefer SVD-based training before deciding whether regularization is necessary. In case SVD is computationally expensive, alternatives for efficient Tikhonov regularization exist, including classical approaches such as Élden’s method [62] and modern techniques for large-scale inverse problems [63].
Our analysis in Section˜V further demonstrated that NGRC achieves superior forecasting performance in hyperparameter regions that balance numerical stability and fitting accuracy. In other words, large values of either the condition number or the closeness of fit tend to degrade performance. This observation suggests another practical heuristic: by monitoring $\kappa(\Psi)$ and the closeness of fit during training, one can readily identify hyperparameter regions unlikely to yield reliable long-term forecasts. When combined with the absence of regularization, this strategy may reduce training computational time.